OpenStudy (anonymous):

Determine whether the sequence could be geometric or arithmetic. If possible, find the common ratio or difference. 1/4, 1/12, 1/36, 1/108

OpenStudy (anonymous):

it is geometric, with a common ratio of 1/3

OpenStudy (anonymous):

1/4 * 1/3 = 1/12 1/12 * 1/3 = 1/36 etc...

OpenStudy (anonymous):

Oh I see. Okay thank you! :) @robz8

OpenStudy (anonymous):

Arithmetic would deal with the common difference. And it would have been like adding or subtracting the same number from each term. Geometric would deal with the common ratio. And it would have been like dividing and multiplying the same number from each term.

OpenStudy (anonymous):

And yes the common ratio would be \[\frac{ 1 }{ 3 }\]

OpenStudy (anonymous):

yw, @some_someone detailed it better than i did